2x+2y−3z = -6
x+6y−4z = -23
3x+y-4z= −2


Found 2 solutions by Alan3354, MathTherapy:
Answer by Alan3354(69443)

2x+2y−3z = -6
x+6y−4z = -23
3x+y-4z= −2
-------------------
Eliminate y terms.
---
 6x + 6y − 9z = -18  Eqn 1 times 3
  x + 6y − 4z = -23
18x + 6y - 24z= −12  Eqn 3 time 6
---
 6x + 6y − 9z = -18  Eqn 1 times 3
  x + 6y − 4z = -23
-------------------------------- Subtract
5x - 5z = 5
==================
  x + 6y − 4z = -23
18x + 6y - 24z= −12  Eqn 3 time 6
------------------------------------Subtract
-17x + 20z = -11
5x - 5z = 5 ---> x = z+5

Sub for x
-17(z-5) + 20z = -11
3z + 85 = -11
3z = -96
z = -32
==================
x = -27
-------
2x+2y−3z = -6
-54 + 2y + 96 = -6
2y = -48
y = -24
--------
(-27,-24,-32)

Answer by MathTherapy(10552)


2x+2y−3z = -6
x+6y−4z = -23
3x+y-4z= −2
No question asked, but anyway: